# Cell-cell communication inference and analysis: biological mechanisms, computational approaches, and future opportunities

### v1

@@ 4.2 Network Methods
The properties of graphs and networks enable intuitive and precise modeling of gene regulatory processes, which is particularly beneficial for tools that integrate downstream signaling and represent genes as nodes. Some tools[RN306, RN302, RN283, RN377SpaTalk] focus on analyzing intracellular signaling while other tools build comprehensive signaling networks that integrate both intercellular and intracellular communication[RN340, RN281, RN401, RN349, sang2025unraveling, RN350, RN454stMLnet]. These tools infer directed regulatory networks from ligands to downstream genes by integrating prior biological knowledge with gene expression data to identify specifically expressed genes or significant signaling pathways, thereby enabling network pruning. For example, to predicts ligand-target links between interacting cell types, NicheNet[RN349, sang2025unraveling] employs the Personalized PageRank (PPR) algorithm to compute a regulatory potential score for all ligand-target gene pairs from prior knowledge sources. This approach prioritizes ligands and their potential target genes by using the ligand of interest as the seed node and calculating a signaling importance score for each gene within the ligand-signaling network. The core approach PPR is formulated as:
@@
To account for spatial constraints of CCC, methods have been developed to emphasis CCC analysis within proximal cells or spots[RN306, RN371Renoir, RN373SpaCeNet, RN377SpaTalk, RN379, RN454stMLnet, RN415FlowSig, RN418, RN409]. For example, SpaTalk[RN377SpaTalk] scores the ligand-receptor-target signaling network between spatially proximal cells by constructing a KNN cell graph network and a prior ligand-receptor-TF knowledge graph and estimating the probability of downstream TF activation via a random walk algorithm. Different from SVCA[RN385SVCA] that models the expression of genes independently of each other, SpaCeNet[RN373SpaCeNet] employs a Gaussian graphical model to decompose the observed cellular profiles into contributions arising from cellular variability and cellular interactions, which captures complex multivariate relationships between genes, and disentangles the intracellular from intercellular gene correlations using spatial conditional independence. Renoir[RN371Renoir] maps spatially resolved ligand-target activities and identifies communication niches by integrating spatial transcriptomics with scRNA-seq data, using a comprehensive neighborhood scoring system that accounts for cellular composition, receptor expression, and spatial proximity to reveal context-specific signaling microenvironments. Building on the earlier method scMLnet[RN350], which constructs a multilayer signaling network comprising three subnetworks including ligand-receptor, receptor-TF and TF-target gene, stMLnet[RN454stMLnet] models the spatial-temporal distribution of ligand concentration $u(x,y,z)$ during the diffusion process using a partial differential equation (PDE), and then computes the signaling strength of ligand-receptor pairs between specific cell pairs. The PDE used to model the ligand diffusion process is formulated as follows:
@@
where $\Delta$ is the Laplace operator; $D$ is the diffusion coefficient; $B_{1}$ represents unit ball indicating the sender cell. Incorporating the law of mass action, the signaling strength $LRS_{j}^{k}$ of the ligand-receptor pair $k$ at the receiver cell $j$ is defined as follows:
@@
The signaling activities $LRS^{1_{t}},LRS^{2_{t}},\dots,LRS^{n_{t}}$ across the receiver cells are used as input of random forest regression model $f_{t}$ to predict the expression of $TG_{t}$.
@@
Unlike methods that rely on prior pathway annotations, CytoTalk[RN318] performs de novo prediction of complete signal transduction pathways mediated by LRIs between two cell types from single-cell transcriptomics. It constructs two intracellular networks based on the mutual information of gene pairs—one for the sender and one for the receiver cells—and connects them via known LRIs. An optimal signaling network that considers genes with high cell type specificity and close connection to high-scoring ligand-receptor pairs are then identified by solving a network propagation-based prize-collecting Steiner forest problem. In addition, LRLoop[RN288] detects feedback signaling loops formed by pairs of LRIs through a network propagation algorithm. Collectively, these methods highlight the increasing diversity and sophistication of modeling strategies tailored to specific biological questions.
@@
Leveraging network-based frameworks, TraSig[RN292] and TICCI[RN434] address intercellular interactions among dynamically transitioning cells during differentiation and development. TraSig first employs a probabilistic graphical model to infer the pseudotemporal ordering of cells, and subsequently applies a sliding window approach to reconstruct gene expression profiles along each trajectory edge, thereby enabling the inference of LRIs between cell clusters that overlap in pseudotime. In contrast, TICCI integrates intercellular association probabilities and communication probabilities to construct a class k-nearest neighbor (KNN) graph with weighted edges, which is then used to assess cellular differentiation states and identify branching trajectories.
@@
For datasets derived from different conditions, network topological metrics can be applied to compare differences between CCC networks[RN280, RN319, RN273, RN397].CINS[RN285] is a multi-step framework that uniquely leverages cell-type proportion changes across conditions in scRNA-seq data to first infer a Bayesian network of differential cell-type interactions and then uses a constrained regression model to identify the key ligand-receptor pairs mechanistically responsible for these interactions, providing a powerful tool for hypothesis generation in case-control studies. Among these, COMUNET[RN280] introduces a dissimilarity metric that accounts not only for the presence or absence of edges but also their weights and directions when comparing edge sets between two layers representing ligand-receptor pairs in a multilayer CCC network. When multiple samples are available, Scriabin[RN352] is a computational framework that performs comparative CCC analysis at true single-cell resolution by constructing cell-cell interaction matrices, employing a binning strategy for large-scale comparisons, and identifying co-expressed interaction programs that are inferred using the Weighted Gene Correlation Network Analysis (WGCNA) framework. Scriabin powerfully reveals heterogeneous and rare communication events obscured by agglomerative methods. MultiNicheNet[RN336] prioritizes differentially expressed and active ligand-receptor pairs across multi-condition and multi-sample datasets by aggregating multiple criteria—such as differential expression, cell type specificity, and sample-level expression—into a final weighted score.
@@ 4.3 Deep Learning Methods
Deep learning has demonstrated remarkable effectiveness across a wide range of applications in the field of single-cell and spatial omics, with a recent increasing efforts of developing computational methods to characterize complex nonlinear relationships in CCC (Figure 3b). These methods excel at learning latent representations of cells and genes, potentially uncovering CCC patterns overlooked by traditional approaches. Particularly notable is the growing application of graph neural networks (GNNs), which naturally model structural relationships and dependencies in graphs representing cellular or gene interactions.
@@
Spatial transcriptomics enables construction of cell-cell spatial graphs that provide contextual information for training neural networks[RN375spaCI, RN363, RN365GraphComm, RN366HoloNet, RN369, RN404CellGAT, zohora2025cellnest, RN422NicheCompass, qi2025interpretable, Xiao20240821608964GITIII]. For instance, CellGAT[RN404CellGAT] employs graph attention networks to integrate scRNA-seq data with protein interaction knowledge, predicting context-specific LRIs with multi-omics evidence. In contrast, DeepCOLOR[RN328DeepCOLOR] uses a variational autoencoder to recover cell-cell colocalization networks at single-cell resolution by integrating single-cell and spatial transcriptomes.
@@
Training neural networks to predict gene expression in receptor cells enables the inference of CCC and the understanding of how LRIs affect downstream targets. HoloNet[RN366HoloNet] and scDCA[RN428scDCA] identify CCC events influencing transcriptional responses through multi-view networks that predict target gene expression and interpret the relative contribution of each view via attention scores. In HoloNet, each view corresponds to a CCC network defined by a specific ligand-receptor pair, whereas in scDCA, each view represents the interaction network between a specific pair of cell types. HoloNet represents the first deep learning method to incorporate detailed signaling mechanisms—including molecular diffusion, cofactors, and LRI-specific interaction ranges—building upon traditional approaches like CellChat. CellAgentChat[RN324CellAgentChat] predicts gene expression profiles from ligand-receptor gene expression and infers regulatory relationships through receptor perturbation analysis. Other tools focus on reconstructing intracellular gene regulatory networks[RN357CLARIFY, RN375spaCI, RN422NicheCompass, RN425PathFinder, RN426]. For example, spaCI[RN375spaCI] maps genes into a latent space using adaptive graph models with attention mechanisms, and constrains interacting gene pairs to be positioned closer together, thereby facilitating the prediction of LRIs and their upstream regulators from the learned representations. CLARIFY[RN357CLARIFY] reconstructs intracellular interaction networks using pre-inferred gene regulatory networks and provides downstream regulatory information for the inference of intercellular communication networks. CLARA[RN406CLARA] is a spatially-aware, transformer-inspired method that infers CCC at the resolution of individual cell pairs by modeling contextual relationships between ligands and receptors within local cellular neighborhoods, providing a highly granular view of the cellular interactome. By integrating gene expression profiles with spatial locations, CCC networks can be directly reconstructed without prior knowledge of ligand-receptor pairs or restriction to immediate neighborhoods[RN362DeepLinc, RN357CLARIFY, RN412DeepTalk, RN433, feng2025orgaccc]. For example, DeepLinc[RN362DeepLinc] can reveal potential long-range interactions by using variational graph autoencoder with adversarial learning. NicheCompass[RN422NicheCompass] interprets cellular interactions through decoder weights, and GITIII[Xiao20240821608964GITIII] characterizes the influence of neighboring cells on the central cell using a single-layer graph transformer model.
@@
To investigate the temporal dynamics of CCC, HiVAE[RN394HiVAE] and CellAgentChat[RN324CellAgentChat] incorporate pseudotime trajectories inferred by existing tools to characterize cellular interactions. HiVAE quantifies information flow between different cell types by computing transfer entropy along the pseudotemporal ordering of cells from scRNA-seq data. CellAgentChat calculates communication scores only between cells within the same pseudotime bin or in adjacent bins, thereby reflecting the temporal proximity of interacting cells. It further trains a neural network to predict gene expression levels of each cell at the next time point, conditioned on the effects of cellular interactions.
@@
Currently, relatively few deep learning methods have been developed to identify differential CCC across biological conditions. Among existing approaches for scRNA-seq data, scTenifoldXct[RN294scTenifoldXct] formulates CCC inference as a manifold alignment problem. It uses neural networks to learn low-dimensional representations of gene expression from two cell types—optionally incorporating known ligand-receptor pairs—and projects a coupled gene-pair similarity matrix from two samples into a unified latent space. Differential LRIs are then identified by computing Euclidean distances between these gene-pair representations. In a different strategy, PathFinder[RN425PathFinder] employs a graph transformer, taking gene expression data and predefined gene-gene interaction paths as input, and is trained to predict the condition label of individual cells. The resulting path weights indicate the relative importance of each interaction in distinguishing between conditions.
@@
Finally, we focus on the technical implementation of deep learning methods, highlighting architectural designs and training strategies. Current approaches employ diverse frameworks—including unsupervised and supervised learning strategies—implemented through variational autoencoders (VAEs)[kingma2022autoencodingvariationalbayes], graph neural networks (GNNs), and Transformers. When applied to spatial transcriptomics, these architectures capture nonlinear dependencies, spatial organization, and interaction ranges inherent to CCC. In unsupervised learning-based tools, most methods employ autoencoders—particularly variational autoencoders (VAEs)[kingma2022autoencodingvariationalbayes]—to learn feature representations that capture potential CCC patterns[RN328DeepCOLOR, RN272SPRUCE, RN394HiVAE, RN422NicheCompass, RN426, RN357CLARIFY, RN362DeepLinc, RN433, feng2025orgaccc, qi2025interpretable]. These models are typically optimized to minimize the discrepancy between the input and reconstructed data, and use the learned latent representations and reconstructed data to infer CCC. For instance, OrgaCCC[feng2025orgaccc] leverages cellular gene expression profiles, cell-cell spatial graph, and known LRIs-derived gene–gene graph to reconstruct both the intercellular interaction network and the LRI network through two orthogonal graph autoencoders operating at the cell and gene levels, respectively. SPRUCE[RN272SPRUCE] captures interaction topics for each cell-cell pair and identify ligand and receptor genes enriched within each topic by employing an embedded topic model built upon a variational autoencoder framework, which achieves unbiased identification of interpretable cell states across multiple datasets by characterizing CCC patterns. Beyond autoencoders, scHyper[RN429scHyper] is a hypergraph neural network-based method that models CCC as a global, high-order network, leveraging the discrepancy between nodes’ intrinsic and contextual embeddings to reconstruct non-linear interaction scores and provide a systems-level view of cellular crosstalk. CellNEST[zohora2025cellnest] adopts graph attention networks (GATs) along with contrastive learning to integrate ligand–receptor information with spatial position at single cell or spot resolution, where unsupervised training is performed by maximizing the Jensen-Shannon divergence between the original network and its corrupted counterpart. Finally, attention scores are used to quantify the probability of CCC, which is computed as follows:
@@
In addition to CellNEST, attention mechanisms have been effectively leveraged by several recent methods such as GITIII[Xiao20240821608964GITIII], AMICI[RN456AMICI] and Steamboat[LiangSteamboat]. GITIII[Xiao20240821608964GITIII] infers spatially-resolved CCC at single-cell resolution from imaging-based spatial transcriptomics data by leveraging a single-layer graph transformer model. It models how a cell’s transcriptional state is shaped by the cell type compositions and expression profiles of its spatial neighbors, providing a powerful solution for datasets with sparse ligand-receptor coverage. Building upon this concept, AMICI[RN456AMICI] utilizes a sparsely regularized, multi-headed attention module to adaptively identify interactions across multiple spatial scales, delineate spatially dependent cellular sub-populations, and connect these interactions to downstream functional consequences in receiver cells. Furthermore, Steamboat[LiangSteamboat] applies a multi-head attention model to dissect cellular interactions across scales by decomposing a cell’s gene expression into distinct components: intrinsic programs, local communication from neighbors, and long-range interactions. A distinctive feature of Steamboat is its capacity for in silico spatial perturbation, enabling the prediction of cellular responses to dynamic microenvironmental changes.
@@
A fundamental challenge in applying supervised learning to CCC inference is the scarcity of gold-standard datasets with experimentally validated interactions, making reliable label definition difficult. To address this, several methods[RN363, RN365GraphComm, RN391, RN284, RN404CellGAT, RN405CellMsg, RN430SEnSCA, peng2025predicting] curate positive labels from experimentally documented LRIs, treating all other pairs as negatives. The trained models are subsequently used to infer previously unannotated LRIs. For example, CellMsg[RN405CellMsg] leverages graph convolutional networks on multimodal protein features to de novo predict a high-confidence ligand-receptor interactome, which it then integrates with scRNA-seq data via a three-point estimation method to quantify communication strength, offering an enhanced, feature-driven approach to decipher CCC networks. Furthermore, SEnSCA[RN430SEnSCA] constructs negative samples of LRIs by replacing the common practice of randomly selecting negative samples from unlabeled ligand-receptor pairs with K-means clustering, thereby reducing label noise and sample distribution bias that could weaken model predictive performance. Other tools adopt alternative strategies, such as thresholding communication scores or aggregating outputs from established statistical methods, to predefine interaction labels as ground truth[RN375spaCI, RN325, RN429scHyper]. Another strategy shifts the prediction target from interactions themselves to downstream effects. In this approach, some models are trained to predict gene expression[RN366HoloNet, RN369, RN324CellAgentChat] or cell type classifications[RN425PathFinder, RN428scDCA], with signaling activity subsequently assessed post-hoc through analysis of model weights or sensitivity analyses.
@@
Recent advances have produced sophisticated hybrid models that integrate multiple deep learning architectures to address the complexity of CCC. Unlike single-architecture approaches, these hybrid frameworks offer modular and flexible designs that can be tailored to specific biological questions. For instance, DeepTalk[RN412DeepTalk] employs a hybrid training strategy combining self-supervised pretraining with supervised fine-tuning to infer CCC at single-cell resolution. The model first learns global connectivity patterns through masked node prediction on large-scale cell graphs, then fine-tunes on specific datasets for binary edge classification, significantly enhancing generalization capability. GraphComm[RN365GraphComm] represents another hybrid approach, constructing directed graphs from scRNA-seq data and applying GATs to integrate protein complex and pathway information. The method leverages a comprehensive knowledge base of over 30,000 protein interaction pairs to capture detailed cellular localization and intracellular signaling patterns, substantially improving CCC inference accuracy. The integration of large language models (LLMs) represents a particularly promising direction. SpaCCC[RN431SpaCCC] fine-tunes pretrained LLMs on gene expression prediction tasks, enabling the embedding of ligands and receptors into a shared latent space. This unified embedding facilitates the identification of potential LRIs and subsequent CCC inference.

### v3

@@ 4.2 Network Methods
The properties of graphs and networks enable intuitive and precise modeling of gene regulatory processes, which is particularly beneficial for tools that integrate downstream signaling and represent genes as nodes. Some tools[118, 92, 18, 153] focus on analyzing intracellular signaling while other tools build comprehensive signaling networks that integrate both intercellular and intracellular communication[35, 70, 126, 24, 149, 33, 191]. These tools infer directed regulatory networks from ligands to downstream genes by integrating prior biological knowledge with gene expression data to identify specifically expressed genes or significant signaling pathways, thereby enabling network pruning. For example, to predict ligand-target links between interacting cell types, NicheNet[24, 149] employs the Personalized PageRank (PPR) algorithm to compute a regulatory potential score for all ligand-target gene pairs from prior knowledge sources. This approach prioritizes ligands and their potential target genes by using the ligand of interest as the seed node and calculating a signaling importance score for each gene within the ligand-signaling network. The core approach PPR is formulated as:
@@
To account for spatial constraints of CCC, methods have been developed to emphasize CCC analysis within proximal cells or spots[118, 146, 151, 153, 175, 191, 4, 36, 180]. For example, SpaTalk[153] scores the ligand-receptor-target signaling network between spatially proximal cells by constructing a KNN cell graph network and a prior ligand-receptor-TF knowledge graph and estimating the probability of downstream TF activation via a random walk algorithm. Different from SVCA[14] that models the expression of genes independently of each other, SpaCeNet[151] employs a Gaussian graphical model to decompose the observed cellular profiles into contributions arising from cellular variability and cellular interactions, which captures complex multivariate relationships between genes, and disentangles the intracellular from intercellular gene correlations using spatial conditional independence. Renoir[146] maps spatially resolved ligand-target activities and identifies communication niches by integrating spatial transcriptomics with scRNA-seq data, using a comprehensive neighborhood scoring system that accounts for cellular composition, receptor expression, and spatial proximity to reveal context-specific signaling microenvironments. Building on the earlier method scMLnet[33], which constructs a multilayer signaling network comprising three subnetworks including ligand-receptor, receptor-TF and TF-target gene, stMLnet[191] models the spatial-temporal distribution of ligand concentration $u(x,y,z)$ during the diffusion process using a partial differential equation (PDE), and then computes the signaling strength of ligand-receptor pairs between specific cell pairs. The PDE used to model the ligand diffusion process is formulated as follows:
@@
where $\Delta$ is the Laplace operator; $D$ is the diffusion coefficient; $B_{1}$ represents a unit ball indicating the sender cell. Incorporating the law of mass action, the signaling strength $LRS_{j}^{k}$ of the ligand-receptor pair $k$ at the receiver cell $j$ is defined as follows:
@@
The signaling activities $LRS^{1_{t}},LRS^{2_{t}},\dots,LRS^{n_{t}}$ across the receiver cells are used as input to random forest regression model $f_{t}$ to predict the expression of $TG_{t}$.
@@
Unlike methods that rely on prior pathway annotations, CytoTalk[79] performs de novo prediction of complete signal transduction pathways mediated by LRIs between two cell types from single-cell transcriptomics. It constructs two intracellular networks based on the mutual information of gene pairs—one for the sender and one for the receiver cells—and connects them via known LRIs. An optimal signaling network that considers genes with high cell type specificity and close connection to high-scoring ligand-receptor pairs is then identified by solving a network propagation-based prize-collecting Steiner forest problem. In addition, LRLoop[189] detects feedback signaling loops formed by pairs of LRIs through a network propagation algorithm. Collectively, these methods highlight the increasing diversity and sophistication of modeling strategies tailored to specific biological questions.
@@
Leveraging network-based frameworks, TraSig[101] and TICCI[58] address intercellular interactions among dynamically transitioning cells during differentiation and development. TraSig first employs a probabilistic graphical model to infer the pseudotemporal ordering of cells, and subsequently applies a sliding window approach to reconstruct gene expression profiles along each trajectory edge, thereby enabling the inference of LRIs between cell clusters that overlap in pseudotime. In contrast, TICCI integrates intercellular association probabilities and communication probabilities to construct a class k-nearest neighbor (KNN) graph with weighted edges, which is then used to assess cellular differentiation states and identify branching trajectories.
@@
For datasets derived from different conditions, network topological metrics can be applied to compare differences between CCC networks[156, 129, 77, 170].CINS[196] is a multi-step framework that uniquely leverages cell-type proportion changes across conditions in scRNA-seq data to first infer a Bayesian network of differential cell-type interactions and then uses a constrained regression model to identify the key ligand-receptor pairs mechanistically responsible for these interactions, providing a powerful tool for hypothesis generation in case-control studies. Among these, COMUNET[156] introduces a dissimilarity metric that accounts not only for the presence or absence of edges but also their weights and directions when comparing edge sets between two layers representing ligand-receptor pairs in a multilayer CCC network. When multiple samples are available, Scriabin[182] is a computational framework that performs comparative CCC analysis at true single-cell resolution by constructing cell-cell interaction matrices, employing a binning strategy for large-scale comparisons, and identifying co-expressed interaction programs that are inferred using the Weighted Gene Correlation Network Analysis (WGCNA) framework. Scriabin powerfully reveals heterogeneous and rare communication events obscured by agglomerative methods. MultiNicheNet[23] prioritizes differentially expressed and active ligand-receptor pairs across multi-condition and multi-sample datasets by aggregating multiple criteria—such as differential expression, cell type specificity, and sample-level expression—into a final weighted score.
@@ 4.3 Deep Learning Methods
Deep learning has demonstrated remarkable effectiveness across a wide range of applications in the field of single-cell and spatial omics, with recent increasing efforts of developing computational methods to characterize complex nonlinear relationships in CCC (Figure 3b). These methods excel at learning latent representations of cells and genes, potentially uncovering CCC patterns overlooked by traditional approaches. Particularly notable is the growing application of graph neural networks (GNNs), which naturally model structural relationships and dependencies in graphs representing cellular or gene interactions.
@@
Spatial transcriptomics enables construction of cell-cell spatial graphs that provide contextual information for training neural networks[163, 197, 155, 103, 56, 200, 212, 19, 144, 187]. For instance, CellGAT[200] employs graph attention networks to integrate scRNA-seq data with protein interaction knowledge, predicting context-specific LRIs with multi-omics evidence. In contrast, DeepCOLOR[97] uses a variational autoencoder to recover cell-cell colocalization networks at single-cell resolution by integrating single-cell and spatial transcriptomes.
@@
Training neural networks to predict gene expression in receptor cells enables the inference of CCC and the understanding of how LRIs affect downstream targets. HoloNet[103] and scDCA[84] identify CCC events influencing transcriptional responses through multi-view networks that predict target gene expression and interpret the relative contribution of each view via attention scores. In HoloNet, each view corresponds to a CCC network defined by a specific ligand-receptor pair, whereas in scDCA, each view represents the interaction network between a specific pair of cell types. HoloNet represents the first deep learning method to incorporate detailed signaling mechanisms—including molecular diffusion, cofactors, and LRI-specific interaction ranges—building upon traditional approaches like CellChat. CellAgentChat[145] predicts gene expression profiles from ligand-receptor gene expression and infers regulatory relationships through receptor perturbation analysis. Other tools focus on reconstructing intracellular gene regulatory networks[17, 163, 19, 54, 173]. For example, spaCI[163] maps genes into a latent space using adaptive graph models with attention mechanisms, and constrains interacting gene pairs to be positioned closer together, thereby facilitating the prediction of LRIs and their upstream regulators from the learned representations. CLARIFY[17] reconstructs intracellular interaction networks using pre-inferred gene regulatory networks and provides downstream regulatory information for the inference of intercellular communication networks. CLARA[174] is a spatially-aware, transformer-inspired method that infers CCC at the resolution of individual cell pairs by modeling contextual relationships between ligands and receptors within local cellular neighborhoods, providing a highly granular view of the cellular interactome. By integrating gene expression profiles with spatial locations, CCC networks can be directly reconstructed without prior knowledge of ligand-receptor pairs or restriction to immediate neighborhoods[105, 17, 193, 199, 55]. For example, DeepLinc[105] can reveal potential long-range interactions by using variational graph autoencoder with adversarial learning. NicheCompass[19] interprets cellular interactions through decoder weights, and GITIII[187] characterizes the influence of neighboring cells on the central cell using a single-layer graph transformer model.
@@
To investigate the temporal dynamics of CCC, HiVAE[112] and CellAgentChat[145] incorporate pseudotime trajectories inferred by existing tools to characterize cellular interactions. HiVAE quantifies information flow between different cell types by computing transfer entropy along the pseudotemporal ordering of cells from scRNA-seq data. CellAgentChat calculates communication scores only between cells within the same pseudotime bin or in adjacent bins, thereby reflecting the temporal proximity of interacting cells. It further trains a neural network to predict gene expression levels of each cell at the next time point, conditioned on the effects of cellular interactions.
@@
Currently, relatively few deep learning methods have been developed to identify differential CCC across biological conditions. Among existing approaches for scRNA-seq data, scTenifoldXct[194] formulates CCC inference as a manifold alignment problem. It uses neural networks to learn low-dimensional representations of gene expression from two cell types—optionally incorporating known ligand-receptor pairs—and projects a coupled gene-pair similarity matrix from two samples into a unified latent space. Differential LRIs are then identified by computing Euclidean distances between these gene-pair representations. In a different strategy, PathFinder[54] employs a graph transformer, taking gene expression data and predefined gene-gene interaction paths as input, and is trained to predict the condition label of individual cells. The resulting path weights indicate the relative importance of each interaction in distinguishing between conditions.
@@
Finally, we focus on the technical implementation of deep learning methods, highlighting architectural designs and training strategies. Current approaches employ diverse frameworks—including unsupervised and supervised learning strategies—implemented through variational autoencoders (VAEs)[96], graph neural networks (GNNs), and Transformers. When applied to spatial transcriptomics, these architectures capture nonlinear dependencies, spatial organization, and interaction ranges inherent to CCC. In unsupervised learning-based tools, most methods employ autoencoders—particularly variational autoencoders (VAEs)[96]—to learn feature representations that capture potential CCC patterns[97, 161, 112, 19, 173, 17, 105, 199, 55, 144]. These models are typically optimized to minimize the discrepancy between the input and reconstructed data, and use the learned latent representations and reconstructed data to infer CCC. For instance, OrgaCCC[55] leverages cellular gene expression profiles, cell-cell spatial graph, and known LRIs-derived gene–gene graph to reconstruct both the intercellular interaction network and the LRI network through two orthogonal graph autoencoders operating at the cell and gene levels, respectively. SPRUCE[161] captures interaction topics for each cell-cell pair and identifies ligand and receptor genes enriched within each topic by employing an embedded topic model built upon a variational autoencoder framework, which achieves unbiased identification of interpretable cell states across multiple datasets by characterizing CCC patterns. Beyond autoencoders, scHyper[106] is a hypergraph neural network-based method that models CCC as a global, high-order network, leveraging the discrepancy between nodes’ intrinsic and contextual embeddings to reconstruct non-linear interaction scores and provide a systems-level view of cellular crosstalk. CellNEST[212] adopts graph attention networks (GATs) along with contrastive learning to integrate ligand–receptor information with spatial position at single cell or spot resolution, where unsupervised training is performed by maximizing the Jensen-Shannon divergence between the original network and its corrupted counterpart. Finally, attention scores are used to quantify the probability of CCC, which is computed as follows:
@@
In addition to CellNEST, attention mechanisms have been effectively leveraged by several recent methods such as GITIII[187], AMICI[75] and Steamboat[108]. GITIII[187] infers spatially-resolved CCC at single-cell resolution from imaging-based spatial transcriptomics data by leveraging a single-layer graph transformer model. It models how a cell’s transcriptional state is shaped by the cell type compositions and expression profiles of its spatial neighbors, providing a powerful solution for datasets with sparse ligand-receptor coverage. Building upon this concept, AMICI[75] utilizes a sparsely regularized, multi-headed attention module to adaptively identify interactions across multiple spatial scales, delineate spatially dependent cellular sub-populations, and connect these interactions to downstream functional consequences in receiver cells. Furthermore, Steamboat[108] applies a multi-head attention model to dissect cellular interactions across scales by decomposing a cell’s gene expression into distinct components: intrinsic programs, local communication from neighbors, and long-range interactions. A distinctive feature of Steamboat is its capacity for in silico spatial perturbation, enabling the prediction of cellular responses to dynamic microenvironmental changes.
@@
A fundamental challenge in applying supervised learning to CCC inference is the scarcity of gold-standard datasets with experimentally validated interactions, making reliable label definition difficult. To address this, several methods[197, 155, 140, 138, 200, 186, 209, 139] curate positive labels from experimentally documented LRIs, treating all other pairs as negatives. The trained models are subsequently used to infer previously unannotated LRIs. For example, CellMsg[186] leverages graph convolutional networks on multimodal protein features to de novo predict a high-confidence ligand-receptor interactome, which it then integrates with scRNA-seq data via a three-point estimation method to quantify communication strength, offering an enhanced, feature-driven approach to decipher CCC networks. Furthermore, SEnSCA[209] constructs negative samples of LRIs by replacing the common practice of randomly selecting negative samples from unlabeled ligand-receptor pairs with K-means clustering, thereby reducing label noise and sample distribution bias that could weaken model predictive performance. Other tools adopt alternative strategies, such as thresholding communication scores or aggregating outputs from established statistical methods, to predefine interaction labels as ground truth[163, 192, 106]. Another strategy shifts the prediction target from interactions themselves to downstream effects. In this approach, some models are trained to predict gene expression[103, 56, 145] or cell type classifications[54, 84], with signaling activity subsequently assessed post-hoc through analysis of model weights or sensitivity analyses.
@@
Recent advances have produced sophisticated hybrid models that integrate multiple deep learning architectures to address the complexity of CCC. Unlike single-architecture approaches, these hybrid frameworks offer modular and flexible designs that can be tailored to specific biological questions. For instance, DeepTalk[193] employs a hybrid training strategy combining self-supervised pretraining with supervised fine-tuning to infer CCC at single-cell resolution. The model first learns global connectivity patterns through masked node prediction on large-scale cell graphs, then fine-tunes on specific datasets for binary edge classification, significantly enhancing generalization capability. GraphComm[155] represents another hybrid approach, constructing directed graphs from scRNA-seq data and applying GATs to integrate protein complex and pathway information. The method leverages a comprehensive knowledge base of over 30,000 protein interaction pairs to capture detailed cellular localization and intracellular signaling patterns, substantially improving CCC inference accuracy. The integration of large language models (LLMs) represents a particularly promising direction. SpaCCC[83] fine-tunes pretrained LLMs on gene expression prediction tasks, enabling the embedding of ligands and receptors into a shared latent space. This unified embedding facilitates the identification of potential LRIs and subsequent CCC inference.